solve 3x/5-2=x/5 Thank you very much
Answer by algebrahouse.com(1079)

3x/5 - 2 = x/5
5[3x/5 - 2 = x/5] {multiplied entire equation by 5 to get rid of fractions}
5(3x/5) - 5(2) = 5(x/5) {multiplied each term by 5}
3x - 10 = x {cancelled and multiplied what's left}
-2x = -10 {subtracted 3x from both sides}
x = 5 {divided both sides by -2}
www.algebrahouse.com